OpenStudy (anonymous):

what is the solution of the proportion x-2/4 = x+10/10?

OpenStudy (campbell_st):

cross multiply to get rid of the denominators 10(x -2) = 4( x + 10) distribute 10x - 20 = 4x + 40 simplifying 6x = 60 you can solve for x

OpenStudy (anonymous):

10